# No-go for the formation of heavy mass Primordial Black Holes in Single Field Inflation

### v5

@@ Abstract
- IX Appendix
@@ I Introduction
The exponential expansion of the nascent universe, dubbed the inflationary paradigm along with the hot big bang model, stands as the most promising theoretical epistemology of the Universe. Invoking inflation from the philosophical rationale to overcome the otherwise muddles like horizon and flatness problems, it was quickly realized that the formation of the structures in our Universe can be a manifestation of the fluctuations generated quantum mechanically during inflation. It was also discovered that large fluctuations associated with specific scales can cause gravitational collapse as the mode reenters the later radiation-dominated era, resulting in the formation of astrophysical objects behaving like black holes and christened to be Primordial Black Holes (PBHs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62] to distinguish them from stellar black holes, which are formed by the death of a star, while keeping Chandrasekhar’s limits in mind. In that regard, the mass of PBHs can be as small as the Planck mass () or the associated cut-off of the effective theory under consideration. PBHs can be the riposte to the curve ball thrown to us by nature: the identity of Dark Matter (DM) [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79]. Apart from the potential solution of the identity conundrum of dark matter, PBHs have gained a lot of interest in the theoretical physics community due to the recent observations of Gravitational Waves (GWs) [80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 77, 93, 78, 79, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105] from merging black holes by Laser Interferometer Gravitational Wave Observatory(LIGO) [106].The large mass and distance of the BHs indicate their primordial origin. In the past few years, the study of PBHs has emerged as one of the most active fields of research in theoretical cosmology. A nice statistical analysis of interest can be found in ref. [54].
@@
The paper is organized as follows: In section II, we discuss the basics of the theoretical background of single field inflationary paradigm, which will going to be extremely useful to understand the rest of the computation performed in this paper. Next, in section III, we provide the detailed computation of tree level scalar power spectrum, where we show the explicitly contributions from SR and USR regions in detail. In both cases, we have studied the behaviour of the tree level scalar power spectrum, which can be analyzed in the sub-horizon (quantum), horizon exit point (semi-classical), and super-horizon (classical) regions, respectively. In section IV, we compute the one-loop corrected unrenormalized scalar power spectrum both from the SR and USR regions. Such one-loop effects are generated in the sub-horizon region, which has a purely quantum mechanical origin and originates from short-range UV modes. We use the cut-off regularization technique to compute the contributions from one-loop integrals, where we introduce both UV and IR cut-offs in the momentum dependent integrals as appearing in both the SR and USR phases. Then, in section V, considering both the contributions from the SR and USR regions, we compute the renormalized version of the one-loop corrected scalar power spectrum by making use of the well-known adiabatic or wave function renormalization scheme [135, 136, 137, 138, 139, 140, 141, 142, 143, 144]. By implementing the renormalization condition correctly within the framework of adiabatic or wave function renormalization scheme, we explicitly compute the expression for the counter terms both in the SR and USR periods, which fix the form of the one-loop renormalized spectrum by removing the quadratic UV divergence. Further using this result we perform the renormalization in the power spectrum which in turn softens the effect of long range modes by shifting the logarithmic IR divergence in the next order of perturbation theory. Next, in section VI, we compute the Dynamical Renormalization Group (DRG) re-summed course grained version of the scalar power spectrum by applying the non-perturbative but convergent exponenciation. Additionally, we discuss the validity of the DRG re-summation method and provide a numerical estimate which justifies its applicability and significance within the present context where we consider sharp transition from SR to USR phase. Using these results, in section VII, we provide a no-go theorem for the mass of the PBH formation in a single field inflationary paradigm. We show that large mass PBHs are not allowed by the present prescription. We also comment on the constraints on the evaporation time scale and on the corresponding number of e-foldings in the USR region using the no-go result. Finally, in section VIII, we conclude with some possible promising future prospects. Additionally, we include two appendices IX.1 and IX.2, where we discuss the details of the computation of the one-loop correction to the primordial power spectrum for the scalar modes using in-in formalism in the USR phase, the technical results of the one-loop momentum integrals in the SR and USR phases, and the appearances of the quadratic UV and logarithmic IR divergent contributions, respectively.
@@ II The single field inflation
where we have introduced a notation which represents the derivative with respect to the conformal time coordinate . Additionally, we have used the definition of the Hubble parameter .
@@ IV Computation of one-loop corrected scalar power spectrum
We now explicitly compute the contribution from the quantum one-loop correction to the power spectrum of the scalar mode during PBH formation from the last term of the third order expanded action as stated in equation(25). For this purpose, we use the well known in-in formalism, which is actually motivated by the Schwinger-Keldysh path integral formalism. Within the framework of in-in formalism, the correlation function of any quantum operator at the fixed conformal time scale can be expressed as:
where and represent the anti-time and time ordering operation in the present context respectively. The interaction Hamiltonian appearing in the above expression can be computed by the following Legendre transformed expression, where we are interested in only on the contribution from last term in the equation(25), which physically represents the leading cubic self-interaction:
The tree level contribution to the two-point correlation function and the corresponding power spectrum is already computed before in this paper. Our job is to fix the non-vanishing contributions from the last three terms in the above given expression at the one-loop level.
@@
For details of the computation, we refer the reader to Appendix IX.1. Here the expressions for the Bogoliubov coefficients and are explicitly written in equation(16) and equation(17). Additionally it is important to note that, we have restricted the momentum integration within a window, by introducing two cut-offs, the IR cut-off and the UV cut-off , to extract the finite contributions from each of the above mentioned integrals. Our further job is to evaluate both of the integrals within the mentioned momentum window. Again it is important to note that, the integral in the sub horizon region and at the horizon crossing scale has the quantum effects and consequently the one-loop contributions become physically meaningful in these above mentioned regions. It is expected that whatever result we derive at the horizon exit point will further propagate to the super horizon region. Now instead of computing two independent integrals we will compute a single momentum integral at the arbitrary conformal time scale and in the final result, we will substitute the values of the conformal time scale and as required by the above mentioned structure of the momentum integrals. To serve the purpose let us compute the following integrals:
where we have considered the contributions which will only give rise to significant cut off dependent divergent contributions in the loop integral. Additionally, two more terms appear which captures the interference between the Bogoliubov coefficients, giving rise to oscillating contribution and no significant divergences. For this reason we have neglected these two contributions. However, to maintain the overall clarity in the computation the appearance of all such contributions in the momentum dependent one-loop contribution in the USR phase are explicitly computed in Appendix IX.2 from which one can clearly observe the corresponding suppression.
@@
Here represents the pivot scale which is expected to be . The last term is the outcome of the sub horizon scale computation. Now in the late time limit, which is actually representing the super-horizon scale (which corresponds to the small momentum modes or commonly identified as the IR modes), the last term vanishes for which we get only logarithmically divergent contribution which will going to survive from this computation at the end and contribute to the one-loop correction to the scalar power spectrum in the SR period. However, since inflation ends at the finite conformal time scale value it is expected to have some small contribution from the UV modes which basically represents the sub horizon fluctuations in the present context. For this reason such UV divergent contributions we have kept intact and in the later part of the paper by doing technical computation we will explicitly show that one can easily remove such contributions from the one-loop corrected power spectrum both from the SR and USR periods of inflation separately by implementing a special renormalization scheme, commonly known as the adiabatic or wave function renormalization technique, which has the inherent power to remove the UV divergences from the theory. However, the IR logarthmic divergences cannot be completely removed after renormalization at the one-loop level. We will further show by implementing the power spectrum renormalization method that in the present context the IR divergent contribution will be shifted to the next-to-next leading order, which is exactly equivalent of performing the computation at the two-loop level. In support of the previously mentioned statement regarding the softening of the IR divergent contribution from the one-loop result is quite understandable from the perspective of the Quantum Field Theory of de Sitter space. It is well known and established fact that de Sitter or its cosmologically relevant version (quasi de Sitter) is IR divergent always and it is almost impossible to remove such divergences from the underlying Quantum Field Theory with the help of presently available tools and techniques. The one-loop correction in the SR phase can be exprerssed by the following expression:
@@
Here we have introduced a Heaviside Theta function at the transition point (SR to USR):
@@ V.1 Wave function/ Adiabatic renormalization method: Complete removal of quadratic UV divergence
where is the characteristic function which is defined for the adiabatic -th order by the following expression:
The above mentioned characteristic function physically represents the adibaticallay regularized conformal time dependent frequency factor appearing at the -th order. Additionally, it is important to mention that to define the adiabatic mode at the -th order in the SR period of inflation for comoving curvature perturbation we have used the Bunch Davies quantum initial condition which fixes the Bogoliubov coefficients at the values as stated in equation (12).
@@
Now in the SR and USR periods of inflation we introduce two counter terms and related to the adiabatic renormalization scheme dependent parameters and by the following expressions:
@@ V.2 Power spectrum renormalization method: Softening of logarithmic IR divergence
Now to softening all the effects coming from one-loop logarithmic IR divergences from SR as well as the USR effects, on SR period, we further define the following renormalized power spectrum for the scalar perturbation:
@@
From the derived structure of the one-loop corrected renormalized power spectrum for the scalar modes one can immediate conclude that no information of the quantum loop correction will be propagated to the pivot scale , where CMB observation takes place. As an immediate consequence in any of the flow of the power spectrum (in the language of quantum field theory, one can say the Renormalization Group (RG) flow equations and the corresponding - functions), which are spectral tilt, running and running of the running of the tilt, are going to be completely independent of the quantum loop effects at the pivot scale . This implies:
These are the outcome of very clever yet extremely logical choice of the counter-term determining renormalization condition. It becomes theoretically justifiable and logically consistent because of the fact that, there should not be any effect appearing from the quantum loops due to having effects from both the SR and USR region, at the pivot scale . Otherwise the loop effects has to be seen and directly tested by the observational probes available till date.
@@ VI Dynamical Renormalization Group (DRG) resummed scalar power spectrum
The final form of the resummed dimensionless power spectrum for the scalar modes can be stated in the following way using the DRG approach 222Note: The uncontrollable increase of the secular terms generally becomes quite troublesome since it invalidates the perturbative computation at the late time scale, especially at the horizon crossing and super-horizon scales. This problem in primordial cosmology can be solved using the DRG approach. :
@@ VII Numerical results and further estimations: A no-go theorem for PBH formation
Finally, it is important to note that during analyzing the representative plots in this paper, we have neglected all other suppressed contributions from the oscillatory terms as well as the inverse power law terms of , as all of them will not be able to significantly change the overall behaviour of the scalar power spectrum in the USR period. For this reason, very small oscillating features do not appear in the final version of the plots. The appearance of all such contributions in the momentum dependent one-loop contribution in the USR phase is explicitly computed in Appendix IX.2 using which one can clearly demonstrate the corresponding suppression.
@@ Acknowledgements
The authors would like to thank Grant J. Mathews and Katherine Freese for bringing our attention to this problem initially. The authors would also like to thank Romesh Kaul, Yogesh and Mohit Kumar Sharma for fruitful discussions. Additionally, the authors of this paper would like to sincerely thank Antonio Riotto, Vincent Vennin, Adam Brown, Maxim Khlopov, Amjad Ashoorioon, Gerasimos Rigopoulos, Ying-li, Shu-Lin Cheng, Da-Shin Lee, Kin-Wang Ng, Jinsu Kim, Theodoros Papanikolaou, Xinpeng Wang, H. V. Ragavendra, Alexandros Karam for fruitful communication. SC would like to thank the work-friendly environment of The Thanu Padmanabhan Centre For Cosmology and Science Popularization (CCSP), SGT University, Gurugram, Delhi-NCR, for providing tremendous support in research and offering the Assistant Professor (Senior Grade) position. SC would also like to thank all the members of our newly formed virtual international non-profit consortium Quantum Aspects of the Space-Time & Matter (QASTM) for elaborative discussions. The work of MRG is supported by DST, Government of India under the Grant Agreement number IF18-PH-228 (DST INSPIRE Faculty Award). The work of MRG and MS is supported by Science and Engineering Research Board (SERB), DST, Government of India under the Grant Agreement number CRG/2022/004120 (Core Research Grant). MS is also partially supported by the Ministry of Education and Science of the Republic of Kazakhstan, Grant No. 0118RK00935 and CAS President’s International Fellowship Initiative (PIFI).
IX Appendix
IX.1 Details of the one loop computation of two point function using in-in formalism in USR phase
We now explicitly compute the contribution from the quantum one-loop correction to the power spectrum of the scalar mode during PBH formation from the last term of the third order expanded action as stated in equation(25). For this purpose we use the well known in-in formalism, which is actually motivated from the Schwinger-Keldysh path integral formalism. Within the framework of in-in formalism, the correlation function of any quantum operator at the fixed conformal time scale can be expressed as:

### v6

@@ Abstract
- A Details of the one loop computation of two point function using in-in formalism in USR phase
- B First step on renormalization
@@ I Introduction
The exponential expansion of the nascent universe, dubbed the inflationary paradigm along with the hot big bang model, stands as the most promising theoretical epistemology of the Universe. Invoking inflation from the philosophical rationale to overcome the otherwise muddles like horizon and flatness problems, it was quickly realized that the formation of the structures in our Universe can be a manifestation of the fluctuations generated quantum mechanically during inflation. It was also discovered that large fluctuations associated with specific scales can cause gravitational collapse as the mode reenters the later radiation-dominated era, resulting in the formation of astrophysical objects behaving like black holes and christened to be Primordial Black Holes (PBHs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76] to distinguish them from stellar black holes, which are formed by the death of a star, while keeping Chandrasekhar’s limits in mind. In that regard, the mass of PBHs can be as small as the Planck mass () or the associated cut-off of the effective theory under consideration. PBHs can be the riposte to the curve ball thrown to us by nature: the identity of Dark Matter (DM) [77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93]. Apart from the potential solution of the identity conundrum of dark matter, PBHs have gained a lot of interest in the theoretical physics community due to the recent observations of Gravitational Waves (GWs) [94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 91, 107, 92, 93, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119] from merging black holes by Laser Interferometer Gravitational Wave Observatory(LIGO) [120].The large mass and distance of the BHs indicate their primordial origin. In the past few years, the study of PBHs has emerged as one of the most active fields of research in theoretical cosmology. A nice statistical analysis of interest can be found in ref. [56].
@@
The paper is organized as follows: In section II, we discuss the basics of the theoretical background of single field inflationary paradigm, which will going to be extremely useful to understand the rest of the computation performed in this paper. Next, in section III, we provide the detailed computation of tree level scalar power spectrum, where we show the explicitly contributions from SR and USR regions in detail. In both cases, we have studied the behaviour of the tree level scalar power spectrum, which can be analyzed in the sub-horizon (quantum), horizon exit point (semi-classical), and super-horizon (classical) regions, respectively. In section IV, we compute the one-loop corrected unrenormalized scalar power spectrum both from the SR and USR regions. Such one-loop effects are generated in the sub-horizon region, which has a purely quantum mechanical origin and originates from short-range UV modes. We use the cut-off regularization technique to compute the contributions from one-loop integrals, where we introduce both UV and IR cut-offs in the momentum dependent integrals as appearing in both the SR and USR phases. Then, in section V, considering both the contributions from the SR and USR regions, we compute the renormalized version of the one-loop corrected scalar power spectrum by making use of the well-known adiabatic or wave function renormalization scheme [149, 150, 151, 152, 153, 154, 155, 156, 157, 158]. By implementing the renormalization condition correctly within the framework of adiabatic or wave function renormalization scheme, we explicitly compute the expression for the counter terms both in the SR and USR periods, which fix the form of the one-loop renormalized spectrum by removing the quadratic UV divergence. Further using this result we perform the renormalization in the power spectrum which in turn softens the effect of long range modes by shifting the logarithmic IR divergence in the next order of perturbation theory. Next, in section VI, we compute the Dynamical Renormalization Group (DRG) re-summed course grained version of the scalar power spectrum by applying the non-perturbative but convergent exponenciation. Additionally, we discuss the validity of the DRG re-summation method and provide a numerical estimate which justifies its applicability and significance within the present context where we consider sharp transition from SR to USR phase. Using these results, in section VII, we provide a no-go theorem for the mass of the PBH formation in a single field inflationary paradigm. We show that large mass PBHs are not allowed by the present prescription. We also comment on the constraints on the evaporation time scale and on the corresponding number of e-foldings in the USR region using the no-go result. Finally, in section VIII, we conclude with some possible promising future prospects. Additionally, we include two appendices A and C, where we discuss the details of the computation of the one-loop correction to the primordial power spectrum for the scalar modes using in-in formalism in the USR phase, the technical results of the one-loop momentum integrals in the SR and USR phases, and the appearances of the quadratic UV and logarithmic IR divergent contributions, respectively.
@@ II The single field inflation
where we have introduced a notation ′ which represents the derivative with respect to the conformal time coordinate . Additionally, we have used the definition of the Hubble parameter .
@@ IV Computation of one-loop corrected scalar power spectrum
We now explicitly compute the contribution from the quantum one-loop correction to the power spectrum of the scalar mode during PBH formation from the last term of the third order expanded action as stated in equation(25). For this purpose, we use the well known in-in formalism, which is actually motivated by the Schwinger-Keldysh path integral formalism. Within the framework of in-in formalism, the two-point correlation function for comoving curvature perturbations at the fixed conformal time scale can be expressed as:
where and represent the anti-time and time ordering operation in the present context respectively. In this expression, for a given mode, represents supper-Hubble radius limit at the end of inflation where the quantum correlation is meaningfully computed in the in-in formalism. The interaction Hamiltonian appearing in the above expression can be computed by the following Legendre transformed expression, where we are interested in only on the contribution from last term in the equation(25), which physically represents the leading cubic self-interaction:
The tree-level contribution to the two-point correlation function and the corresponding power spectrum have already been computed in this paper. Our job is to fix the non-vanishing contributions from the last three terms in the above-given expression at the one-loop level as appearing in equation (IV). The meaning of each of the terms as appealing in equation (IV) is further discussed in detail in the Appendix A.
@@
Here represents the tree-level contribution to the power spectrum as appearing in the SR period. Further, it is important to note that, the symbols and characterizing the values of the second slow-roll parameters at the SR to USR transition time scale and at the end of USR as well as the end of inflation scale respectively. Here is the most important scale where sharp transition is implemented to describe SR to USR phase change. We refer the reader to Appendix A for details. Here the expressions for the Bogoliubov coefficients and are explicitly written in equation(16) and equation(17). Additionally it is important to note that, we have restricted the momentum integration within a window, by introducing two cut-offs, the Infra-Red (IR) cut-off and the Ultra-Violate (UV) cut-off , to extract the finite contributions from each of the above mentioned integrals. Our further job is to evaluate both of the integrals within the mentioned momentum window. Again it is important to note that, the integral in the sub horizon region and at the horizon crossing scale has the quantum effects and consequently the one-loop contributions become physically meaningful in these above mentioned regions. It is expected that whatever result we derive at the horizon exit point will further propagate to the super horizon region. Now instead of computing two independent integrals we will compute a single momentum integral at the arbitrary conformal time scale and in the final result, we will substitute the values of the conformal time scale and as required by the above mentioned structure of the momentum integrals. To serve the purpose let us compute the following integrals:
where we have considered the contributions which will only give rise to significant cut off dependent divergent contributions in the loop integral. Additionally, two more terms appear which captures the interference between the Bogoliubov coefficients, giving rise to oscillating contribution and no significant divergences. For this reason we have neglected these two contributions. However, to maintain the overall clarity in the computation the appearance of all such contributions in the momentum dependent one-loop contribution in the USR phase are explicitly computed in Appendix C from which one can clearly observe the corresponding suppression.
@@
Here represents the pivot scale which is expected to be . The last term is the outcome of the sub horizon scale computation. Now in the limit, which is actually representing the super-horizon scale (which corresponds to the small momentum modes or commonly identified as the IR modes), the last term vanishes for which we get only logarithmically divergent contribution which will going to survive from this computation at the end and contribute to the one-loop correction to the scalar power spectrum in the SR period. However, since inflation ends at the finite conformal time scale value it is expected to have some small contribution from the UV modes which basically represents the sub horizon fluctuations in the present context. For this reason such UV divergent contributions we have kept intact and in the later part of the paper by doing technical computation we will explicitly show that one can easily remove such contributions from the one-loop corrected power spectrum both from the SR and USR periods of inflation separately by implementing a special renormalization scheme, commonly known as the adiabatic or wave function renormalization technique, which has the inherent power to remove the UV divergences from the theory. However, the IR logarthmic divergences cannot be completely removed after renormalization at the one-loop level. We will further show by implementing the power spectrum renormalization method that in the present context the IR divergent contribution will be shifted to the next-to-next leading order, which is exactly equivalent of performing the computation at the two-loop level. In support of the previously mentioned statement regarding the softening of the IR divergent contribution from the one-loop result is quite understandable from the perspective of the Quantum Field Theory of de Sitter space. It is well known and established fact that de Sitter or its cosmologically relevant version (quasi de Sitter) is IR divergent always and it is almost impossible to remove such divergences from the underlying Quantum Field Theory with the help of presently available tools and techniques. The one-loop correction in the SR phase can be exprerssed by the following expression:
@@
Here we have introduced a Heaviside Theta function at the transition point (SR to USR) to describe the sharp transition:
@@ V.1 Wave function/ Adiabatic renormalization method: Complete removal of quadratic UV divergence
where is the characteristic function which is defined for the adiabatic -th order by the following expression [150]:
Here ′ denotes the conformal time derive. The above mentioned characteristic function physically represents the adibaticallay regularized conformal time dependent frequency factor appearing at the -th order. Additionally, it is important to mention that to define the adiabatic mode at the -th order in the SR period of inflation for comoving curvature perturbation we have used the Bunch Davies quantum initial condition which fixes the Bogoliubov coefficients at the values as stated in equation (12).
@@
Now in the SR and USR periods of inflation we introduce two counter terms and related to the adiabatic renormalization scheme dependent parameters and by the following expressions222 Here refer to the points of subtraction.:
@@ V.2 Power spectrum renormalization method: Softening of logarithmic IR divergence
Now to softening all the effects coming from one-loop logarithmic IR divergences from SR as well as the USR effects, on SR period, we further define the following renormalized power spectrum for the scalar perturbation 333See the appendix B for the more details on the justification of using two renormalization schemes under a single theoretical framework.:
@@
From the derived structure of the one-loop corrected renormalized power spectrum for the scalar modes one can immediate conclude that no information of the quantum loop correction will be propagated to the pivot scale , where CMB observation takes place. As an immediate consequence in any of the flow of the power spectrum (in the language of quantum field theory, one can say the Renormalization Group (RG) flow equations and the corresponding - functions), which are spectral tilt, running and running of the running of the tilt, are going to be completely independent of the quantum loop effects at the pivot scale . This implies the constraints on the following cosmological - functions:
These are the outcome of very clever yet extremely logical choice of the counter-term determining renormalization condition. It becomes theoretically justifiable and logically consistent because of the fact that, there should not be any effect appearing from the quantum loops due to having effects from both the SR and USR region, at the pivot scale . Otherwise, the loop effects has to be seen and directly tested by the observational probes available till date.
@@ VI Dynamical Renormalization Group (DRG) resummed scalar power spectrum
The final form of the resummed dimensionless power spectrum for the scalar modes can be stated in the following way using the DRG approach 444Note: The uncontrollable increase of the secular terms generally becomes quite troublesome since it invalidates the perturbative computation at the late time scale, especially at the horizon crossing and super-horizon scales. This problem in primordial cosmology can be solved using the DRG approach. :
@@ VII Numerical results and further estimations: A no-go theorem for PBH formation
Finally, it is important to note that during analyzing the representative plots in this paper, we have neglected all other suppressed contributions from the oscillatory terms as well as the inverse power law terms of , as all of them will not be able to significantly change the overall behaviour of the scalar power spectrum in the USR period. For this reason, very small oscillating features do not appear in the final version of the plots. The appearance of all such contributions in the momentum dependent one-loop contribution in the USR phase is explicitly computed in Appendix C using which one can clearly demonstrate the corresponding suppression.
@@ Acknowledgements
The authors would like to thank Grant J. Mathews and Katherine Freese for bringing our attention to this problem initially. The authors would also like to thank Romesh Kaul, Yogesh and Mohit Kumar Sharma for fruitful discussions. Additionally, the authors of this paper would like to sincerely thank Antonio Riotto, Vincent Vennin, Adam Brown, Maxim Khlopov, Amjad Ashoorioon, Gerasimos Rigopoulos, Ying-li, Shu-Lin Cheng, Da-Shin Lee, Kin-Wang Ng, Jinsu Kim, Theodoros Papanikolaou, Xinpeng Wang, H. V. Ragavendra, Alexandros Karam for fruitful communication. SC would like to thank the work-friendly environment of The Thanu Padmanabhan Centre For Cosmology and Science Popularization (CCSP), SGT University, Gurugram, Delhi-NCR, for providing tremendous support in research and offering the Assistant Professor (Senior Grade) position. SC would like to thank The North American Nanohertz Observatory for Gravitational Waves (NANOGrav) collaboration and the National Academy of Sciences (NASI), Prayagraj, India, for being elected as an associate member and the member of the academy respectively. SC would also like to thank all the members of our newly formed virtual international non-profit consortium Quantum Aspects of the Space-Time & Matter (QASTM) for elaborative discussions. The work of MRG is supported by DST, Government of India under the Grant Agreement number IF18-PH-228 (DST INSPIRE Faculty Award). The work of MRG and MS is supported by Science and Engineering Research Board (SERB), DST, Government of India under the Grant Agreement number CRG/2022/004120 (Core Research Grant). MS is also partially supported by the Ministry of Education and Science of the Republic of Kazakhstan, Grant No. 0118RK00935 and CAS President’s International Fellowship Initiative (PIFI).
Appendix A Details of the one loop computation of two point function using in-in formalism in USR phase
We now explicitly compute the contribution from the quantum one-loop correction to the power spectrum of the scalar mode during PBH formation from the last term of the third order expanded action as stated in equation(25). For this purpose we use the well known in-in formalism, which is actually motivated from the Schwinger-Keldysh path integral formalism. Within the framework of in-in formalism, the correlation function of any quantum operator at the fixed conformal time scale can be expressed as [51]: